# Adaptive Reduced Multilevel Splitting

### v1

@@ 4.4 AMS-based importance sampling estimator
Using the sequence of empirical distributions $\{\mu_{{l^{(k)}}}^{{N,(k)}}\}_{k=1}^{K}$ generated by AMS simulations, we can write a closed-form approximation of the importance sampling estimator $\hat{\gamma}_{l_{\max}}^{H}$ of $\gamma^{\ast}_{l_{\mathrm{max}}}$ given in (3.1). This AMS approximation of estimator $\hat{\gamma}_{l_{\max}}^{H}$ is given by
$\hat{\gamma}^{H,N}_{l_{\max}}(\varphi)\displaystyle\mathop{=}^{{\rm def}}\frac{1}{H}\sum_{k=J+1}^{J+H}\frac{\mu_{{l^{(k)}}}^{{N,(k)}}(\mathrm{e}^{\tau^{(k)}E^{(k)}})}{\mathrm{e}^{\tau^{(k)}E^{(k)}(X_{k+1})}}{Z_{l^{(k)}}^{N,(k)}}\varphi(X_{k+1}){\,{\rm l}{\rm l}_{S^{\ast}>l_{\max}}(X_{k+1})},$ (4.10)
In particular, the AMS approximation of the probability estimate $\hat{p}_{l_{\max}}$ of $p^{\ast}_{l_{\max}}$ given in (3.1) is defined as
$\hat{p}^{H,N}_{l_{\max}}\displaystyle\mathop{=}^{{\rm def}}\hat{\gamma}^{H,N}_{l_{\max}}(\,{\rm l}{\rm l}),$ (4.11)
AMS also provides as by-products a sequence of estimator of $\pi(\{S^{(k)}\geq{l_{\max}}\})$ given by
$\hat{p}^{N}_{l_{\max}}\displaystyle\mathop{=}^{{\rm def}}{Z_{l^{(k)}}^{N,(k)}}\mu_{l^{(k)}}^{N,(k)}(\,{\rm l}{\rm l}_{S^{(k)}>{l_{\max}}}),$ (4.12)
that can be an increasingly accurate approximation of $p^{\ast}_{l_{\max}}$ in the case where the reduced score $S^{(k)}$ converges to $S^{\ast}$ as $k$ increases.
@@ 4.5 The ARMS algorithm
The complete methodology proposed in this work is taken up by the algorithm presented below named adaptive reduced multilevel splitting (ARMS). ARMS is a practical implementation of the idealized importance sampling estimator (Algorithm 1), built on the AMS simulation framework described above.
@@
1:level of interest $l_{\mathrm{max}}$, initial reduced score $S^{(1)}$ and error $E^{(1)}$, budget of snapshots $K$, the quantile $\theta\in(0,1)$ proposing the bridge level and level increases in AMS, number of hits of rare event target $J$, number of particles $N$ for AMS, worst-case log cost threshold $e$, reduced score precision $\epsilon\ll e$, boolean $restart$ variable for restarting or bridging AMSs. $\triangleright$ ( initialization)
2:Set $l^{(0)}=l_{\mathrm{bridge}}(1)=-\infty$, $H=0$, $j=0$, $k=0$, $\tau=+\infty$, $Z_{l^{(0)}}^{N,(0)}=1$
3:Simulate $\mu_{l_{\mathrm{bridge}}(1)}^{N,(1)}$ by sampling distribution $\pi$
@@
6:     Use Algorithm 2 (see Section 4.2) from $l_{0}=l_{\mathrm{bridge}}(k)$ to:
  • $\bullet$ estimate the critical level $l^{(k)}\in[l_{\mathrm{bridge}}(k),l_{\mathrm{max}}]$
  • $\bullet$ simulate $\mu_{l^{(k)}}^{N,(k)}$ starting from $\mu_{l_{\mathrm{bridge}}(k)}^{N,(k)}$
@@
7:    Let $(\Xi_{1},\ldots,\Xi_{N})$ be particles in ${\mu^{N,{(k)}}_{l^{(k)}}}$ and draw
$X\sim\frac{\sum_{n=1}^{N}e^{\tau E^{(k)}}\delta_{\Xi_{n}}}{\sum_{n=1}^{N}e^{\tau E^{(k)}({\Xi_{n}})}}$
@@
15:    if $j=J$ then set the learning function parameter as $\tau=0$ end if $\triangleright$ ( importance sampling)
16:    if  $j>J$ then
@@
$\hat{\gamma}^{H,N}_{l_{\max}}(\varphi)\leftarrow\frac{(H-1)\hat{\gamma}^{H-1,N}_{l_{\max}}(\varphi)+Z^{N,{(k)}}_{l^{(k)}}\varphi(X)\,{\rm l}{\rm l}_{S^{\ast}(X)>l_{\max}}}{H}$
@@
  • $\bullet$ find the largest feasible pair $(k_{\mathrm{bridge}},l_{\mathrm{bridge}}(k+1))$
  • $\bullet$ simulate $\mu^{N,(k+1)}_{l_{\mathrm{bridge}}(k+1)}$ from $\mu^{N,(k_{\mathrm{bridge}})}_{l^{(k_{\mathrm{bridge}})}}$
  • $\bullet$ compute the associated normalisation $Z^{N,(k+1)}_{l_{\mathrm{bridge}}(k+1)}$.
24:         elsenormal-$\triangleright$ ( or AMS reinitialization)
25:             Set $l_{\mathrm{bridge}}(k+1)=-\infty$
26:             Set ${Z^{N,{(k+1)}}_{l_{\mathrm{bridge}}(k+1)}}=1$
27:             Simulate $\mu_{l_{\mathrm{bridge}}(k+1)}^{N,(k+1)}$ by sampling distribution $\pi$
@@
30:         Set $l_{\mathrm{bridge}}(k+1)=l^{(k)}$
31:         Set $Z^{N,(k+1)}_{l_{\mathrm{bridge}}(k+1)}=Z^{N,(k)}_{l^{(k)}}$, $\mu^{N,(k+1)}_{l_{\mathrm{bridge}}(k+1)}=\mu^{N,(k)}_{l^{(k)}}$
@@
34:Return Estimate $\hat{\gamma}^{H,N}_{l_{\max}}(\varphi)$
@@
The first block consists in increasing the level until reaching the first non-achievable level presented in Section 2.2 and implemented in practice by an AMS routine in Section 4.2. In short, it simulates the distribution flow $l\mapsto\mu^{N,(k)}_{l}$ with $l\in[{l_{\mathrm{bridge}}(k)},l^{(k)}]$, where $l^{(k)}$ is the critical level achievable for the current reduced model $S^{(k)}$.
The second block samples a new sample according to the proposal $\mu^{N,(k)}_{l^{(k)}}$, penalized by some learning function. It then computes the related snapshot and updates the reduced score. This block follows the guideline presented in Section 2.3.
The third block performs importance sampling with the proposal $\mu^{N,(k)}_{l^{(k)}}$ and updates the IS estimate (4.10). The learning function is also updated in this block.
The fourth and final block is the restart or bridge routine presented in Section 2.4 and implemented in practice by an AMS step in Section 4.3. In short, if the restart option is untriggered ($restart=false$), then the bridging procedure searches in past simulations to find a certain pair $(k_{\mathrm{bridge}},l_{\mathrm{bridge}}(k+1))$ which enables to simulate from $\mu^{N,(k_{\mathrm{bridge}})}_{l^{(k_{\mathrm{bridge}})}}$ towards $\mu^{N,(k+1)}_{l_{\mathrm{bridge}}(k+1)}$ in a single AMS step. Otherwise (in the case $restart=true$), a fresh AMS simulation is initialized at each score approximation update. We mention that the bridging routine is avoided in the case where the stopping criterion (2.8) is reached. Indeed, in this case, no bridging is needed, since the reduced score is not updated anymore and the next initial level can simply be set to $l^{(k)}$.
@@ 5 Numerical Simulations
We first consider an analytical toy model approximated by cubic spline interpolation. The rare event simulation is in this case one-dimensional but nonetheless challenging, as the score presents several maxima and moreover splines may produce highly oscillating approximations. We then focus our attention on a more realistic scenario. The rare event is defined through the high-dimensional solution of an elliptic PDE with random inputs. The PDE solution is approximated by a popular reduced basis approach.
5.1 Two examples
@@
Fig. 1: Illustration of the algorithm adaptive behavior (example 1). The four plots illustrate the algorithm at iteration $k=1,20,40,60$. They display in blue the score function $S^{\ast}$ (the rare event being $S^{\ast}(X)\geq 90$), in red the reduced score approximation $S^{(k)}$, in green the snapshots $(X_{1},\ldots,X_{k})$ used to build the reduced score.
@@ 5.1.1 Example #1: spline approximation of a closed-form score
In a first numerical experiment, we design a one-dimensional analytical toy model. Cubic spline data interpolation is used to build the surrogate $S(x)$ from the snapshots and the error bound is set to the absolute value of the true error $E(x)=|S(x)-S^{\ast}(x)|$.
Using spline interpolation is not recommended in practice, as it easily creates wide oscillations that can artificially create spurious high score areas that misleads the sampling routines. It is nonetheless an interesting benchmark for difficult situations where the surrogate modeling is sub-optimal.
The score function, depicted in Figure 1, is defined for $x\in\mathcal{X}=\mathbb{R}_{+}$ by
$S^{\ast}(x)=\left\{\begin{aligned} l_{\max}\quad&\textrm{if}\quad x\leq 1/l_{\max}\\ 1/x+f(x)\quad&\textrm{else}\end{aligned}\right.,$
where $f(x)=\alpha\left(\,{\rm l}{\rm l}_{x_{0}\leq x<x_{1}}\sin(x-x_{0})^{2}+\,{\rm l}{\rm l}_{x_{1}\leq x}(\sin(x_{1}-x_{0})^{2}-\beta(x-x_{1}))\right),$ given parameters $x_{1}>x_{0}>1/l_{\max}$ and $\alpha,\beta>0$. The reference distribution $\pi$ is a log-normal distribution such that $\ln(x)\sim\mathcal{N}(1.5,1.5)$. The parameter are set to $x_{0}=0.5$, $x_{1}=5$, $\alpha=15$, $\beta=0.1$ and $l_{\max}=90$, corresponding to the rare event probability (computed analytically) $p^{*}_{l_{\max}}=\pi(\,{\rm l}{\rm l}_{S^{\ast}>l_{\max}})=2.18e-8$.
@@ 5.1.2 Example #2: reduced basis approximation of a PDE score
In our second numerical experiment, the score function is defined as $S^{\ast}(x)=\|\Psi^{\ast}(x)\|_{\infty}$ with $\Psi^{\ast}(x)$ being the high-fidelity numerical approximation of the solution of a bi-dimensional PDE parametrized by a $4$-dimensional random vector $x=(x_{1}\cdots x_{4})^{\intercal}\in\mathcal{X}=\mathbb{R}_{+}^{4}$. The chosen PDE corresponds to the well-studied thermal block problem, which models heat diffusion in an heterogeneous media. The reference distribution for $x$ will be a log-normal distribution with independent components with $\ln(x_{q})\sim\mathcal{N}(1.5,1.5)$, for $q=1,\ldots,4$. The probability of interest will correspond to the occurrence of a temperature exceeding a given threshold $l_{\max}=3$ in some point of the spatial domain $\Omega=[0,1]^{2}$. The PDE solution will be approximated using the popular reduced basis (RB) method.
@@
where the function $\kappa(\cdot;x):\Omega\to\mathbb{R}_{+}$ represents a parametric diffusion coefficient function.
The high-fidelity approximation $\Psi^{\ast}(x)$ of the solution of this diffusion problem entails the solution of a variational problem. Precisely, let $\mathcal{V}$ denote the Hilbert space $\mathcal{H}^{1}_{0}(\Omega)$ and $\|\cdot\|_{\mathcal{V}}$ the norm induced by the inner product defined over $\mathcal{V}$. For any $x\in\mathcal{X}$, we define the solution $\Psi^{\star}(x)\in\mathcal{V}$ of the variational problem $a(\Psi(x),\nu;x)=\int_{\Omega}\nu\mathrm{d}\Omega,\quad\forall\nu\in\mathcal{V},$ where $a:\mathcal{V}\times\mathcal{V}\to\mathbb{R}$ is the bilinear form $a(u,\nu;x)=\int_{\Omega}\kappa(\cdot;x)\nabla u\cdot\nabla\nu d\Omega$. We consider the piecewise constant diffusion coefficient $\kappa(\cdot;x)=\sum_{q=1}^{4}\,{\rm l}{\rm l}_{z\in\Omega_{q}}(\cdot)x_{q}$ with the partition $\Omega=\cup_{q=1}^{4}\Omega_{q}$ and $\Omega_{q}$’s being non-overlapping squares. Setting the parameter space $\mathcal{X}=(\mathbb{R}_{+})^{4}$ implies that the bilinear form is strongly coercive, and the Lax-Milgram lemma ensures the variational problem has a unique solution.
@@
with the energy norm induced by the bilinear form $\|\nu\|=\sqrt{a(\nu,\nu;x)}$. The high-fidelity approximation $\Psi^{\ast}(x)$ is obtained in practice by solving a large system of $h$ linear equations. [21]

### v3

@@ 3.4 AMS-based importance sampling estimator
Using the sequence of empirical distributions $\{\mu_{{l^{(k)}}}^{{N,(k)}}\}_{k=1}^{K}$ generated by AMS simulations, we can write a closed-form approximation of the importance sampling estimator $\hat{\gamma}_{l_{\max}}^{H}$ of $\gamma^{\ast}_{l_{\mathrm{max}}}$ defined by (2.9). The AMS approximation of the estimator $\hat{\gamma}_{l_{\max}}^{H}$ is defined by
$\hat{\gamma}^{H,N}_{l_{\max}}(\varphi)\displaystyle\mathop{=}^{{\rm def}}\frac{1}{H}\sum_{k=K_{j_{0}}+1}^{K_{j_{0}}+H}{Z_{l^{(k)}}^{N,(k)}}\varphi(X_{k+1}){\,{\rm l}{\rm l}_{S^{\ast}>l_{\max}}(X_{k+1})},$ (3.10)
In particular, the AMS approximation of the estimator $\hat{\gamma}^{H}_{l_{\max}}(\,{\rm l}{\rm l})$ of $p^{\ast}=\gamma^{\ast}_{l_{\max}}(\,{\rm l}{\rm l})$ is
$\hat{p}_{IS}\displaystyle\mathop{=}^{{\rm def}}\hat{\gamma}^{H,N}_{l_{\max}}(\,{\rm l}{\rm l}),$ (3.11)
AMS also provides as by-products a sequence of approximate estimators of $\pi(\{S^{(k)}\geq{l_{\max}}\})$ given by
$\hat{p}_{AMS}\displaystyle\mathop{=}^{{\rm def}}{Z_{l^{(k)}}^{N,(k)}}\mu_{l^{(k)}}^{N,(k)}(\,{\rm l}{\rm l}_{S^{(k)}>{l_{\max}}}),$ (3.12)
This estimator can be an increasingly accurate approximation of $p^{\ast}$, but as our numerical experiments show, it can also be highly biased, even in the case where the reduced score $S^{(k)}$ converges uniformly to $S^{\ast}$ as $k$ increases.
@@ 3.5 The ARMS algorithm
The complete methodology proposed in this work is summarized in the complete algorithm presented below named adaptive reduced multilevel splitting (ARMS). ARMS is a practical implementation of the idealized importance sampling estimator (Algorithm 1), built on the AMS simulation framework described above.
@@
1:level of interest $l_{\mathrm{max}}$, initial reduced score $S^{(1)}$ and error $E^{(1)}$, budget of snapshots $K$, the quantile $\theta\in(0,1)$ proposing the bridge level and level increases in AMS, number of hits of rare event target $j_{0}$, number of particles $N$ for AMS, worst-case log cost threshold $c$, reduced score precision $\epsilon\ll e$, boolean $restart$ variable for restarting or bridging after a reduced model update, initial learning parameter $\tau_{0}$. $\triangleright$ ( initialization)
2:Set $l^{(0)}=l_{b}(1)=-\infty$, $H=0$, $j=0$, $k=0$, $\tau=\tau_{0}$, $Z_{l^{(0)}}^{N,(0)}=1$
3:Set $\mu_{l_{b}(1)}^{N,(1)}$ by drawing a $N$-sample of the reference distribution $\pi$
@@
6:     Use Algorithm 2 (see Section 3.2) from $l_{b}=l_{b}(k)$ to:
  • $\bullet$ estimate the critical level $l^{(k)}\in[l_{b}(k),l_{\mathrm{max}}]$
  • $\bullet$ simulate $\mu_{l^{(k)}}^{N,(k)}$ starting from $\mu_{l_{b}(k)}^{N,(k)}$
@@
7:    Draw
$X\sim\frac{{\mu^{N,{(k)}}_{l^{(k)}}}e^{\tau E^{(k)}}}{{\mu^{N,{(k)}}_{l^{(k)}}}(e^{\tau E^{(k)}})}$
@@
15:    if $j<j_{0}$ then update the learning function parameter $\tau$ else set $\tau=0$ end if $\triangleright$ ( importance sampling)
16:    if  $j>j_{0}$ then
@@
$\hat{\gamma}^{H,N}_{l_{\max}}\leftarrow\frac{1}{H}\left((H-1)\hat{\gamma}^{H-1,N}_{l_{\max}}+Z^{N,{(k)}}_{l^{(k)}}\,{\rm l}{\rm l}_{S^{\ast}(X)>l_{\max}}\delta_{X}\right)$
@@
  • $\bullet$ find the largest feasible pair $(k_{b},l_{b}(k+1))$
  • $\bullet$ simulate $\mu^{N,(k+1)}_{l_{b}(k+1)}$ from $\mu^{N,(k_{b})}_{l^{(k_{b})}}$
  • $\bullet$ compute the associated normalisation $Z^{N,(k+1)}_{l_{b}(k+1)}$.
24:         $\triangleright$ ( or AMS reinitialization)
25:             Set $l_{b}(k+1)=-\infty$
26:             Set ${Z^{N,{(k+1)}}_{l_{b}(k+1)}}=1$
27:             Set $\mu_{l_{b}(k+1)}^{N,(k+1)}$ by drawing a $N$-sample of the reference distribution $\pi$
@@
30:         Set $l_{b}(k+1)=l^{(k)}$
31:         Set $Z^{N,(k+1)}_{l_{b}(k+1)}=Z^{N,(k)}_{l^{(k)}}$, $\mu^{N,(k+1)}_{l_{b}(k+1)}=\mu^{N,(k)}_{l^{(k)}}$
@@
34:Return Estimate $\hat{\gamma}^{H,N}_{l_{\max}}$
@@
The first block (lines 4 to 5) consists in increasing the level until reaching the first non-achievable level presented in Section 2.2 and implemented in practice by an AMS routine in Section 3.2. In short, it simulates the distribution flow $l\mapsto\mu^{N,(k)}_{l}$ with $l\in[{l_{b}(k)},l^{(k)}]$, where $l^{(k)}$ is the critical level where IS of the true target $\eta^{\ast}_{l^{(k)}}$ by the current reduced model $S^{(k)}$ is expected to be too expensive.
The second block (lines 6 to 14) samples a new sample according to the proposal $\mu^{N,(k)}_{l^{(k)}}$, penalized by some learning function in the early iterations. It then computes the related snapshot, updates the reduced score and finally updates the learning function. This block follows the guideline presented in Section 2.3.
The third block (lines 15 to 19) is the estimation phase beginning when the algorithm has reached the rare event in some sense. This block performs importance sampling with the proposal $\mu^{N,(k)}_{l^{(k)}}$ and updates the IS estimate (3.10).
The fourth and final block (lines 20 to 31) is the restart or bridge routine presented in Section 2.4 and implemented in practice by an AMS step in Section 3.3. In short, if the restart option is untriggered ($restart=false$), then the bridging procedure searches in past simulations to find a certain pair $(k_{b},l_{b}(k+1))$ which enables to simulate from $\mu^{N,(k_{b})}_{l^{(k_{b})}}$ towards $\mu^{N,(k+1)}_{l_{b}(k+1)}$ in a single AMS step. Otherwise (in the case $restart=true$), a fresh AMS simulation is initialized at each score approximation update. We mention that the bridging routine is avoided in the case where the stopping criterion (2.8) is reached. Indeed, in this case, no bridging is needed, since the reduced score is not updated anymore and the next initial level can simply be set to $l^{(k)}$.
@@ 4 Numerical Simulations
We first consider a toy example given by an explicit one dimensional true score. Reduced modeling is performed by cubic spline interpolation. The rare event simulation is in this case one-dimensional but nonetheless challenging, as the score presents several maxima and moreover splines may produce highly oscillating approximations. We then focus our attention on a more realistic scenario in which the rare event is defined through the high-dimensional solution of an elliptic PDE with random inputs. The PDE solution is approximated by a reduced basis approach which yields the reduced models. All our numerical simulations are carried out using an implementation of the ARMS algorithm in the Julia language, available at https://gitlab.inria.fr/pheas/arms.
4.1 Examples
@@
Figure 1: Illustration of the algorithm adaptive behavior (example #1). The four plots illustrate the algorithm at iteration $k=1,20,40,60$. They display in blue the score function $S^{\ast}$ called full order model (FOM) (the rare event being $S^{\ast}(X)\geq 90$), in red the reduced score approximation $S^{(k)}$ called reduced order model (ROM), in green the snapshots $(X_{1},\ldots,X_{k})$ used to build the reduced score.
We will consider the problem of estimation of $p^{\ast}=\pi(\,{\rm l}{\rm l}_{S^{\ast}>l_{\max}})$ with $S^{\ast}(X)=\mathrm{score}(\Psi^{\ast}(X))$, for various models $\Psi^{\ast}$ and reduced models $\Psi$, various dimensions $q=\dim(X)$ of the parameter space, various model dimensions $h=\dim(\Psi^{\ast}(X))$ and different score functions.
$\Psi^{\ast}$ $\Psi$ $\dim(X)$ $\dim(\Psi^{\ast}(X))$ $\mathrm{score}$ $p^{\ast}$ $l_{\max}$
example #1 analytical spline 1 1 ||$\cdot$|| ⋅ | 2.18e-08 90
example #2a PDE RB 4 5101 $L^{1}$ 6.31e-04 0.5
example #2b PDE RB 4 5101 $L^{\infty}$ 1.47e-04 3.0
example #2c PDE RB 25 5101 $L^{1}$ 2.48e-14 0.5
We detail hereafter these examples.
@@ 4.1.1 Example #1: spline approximation of a closed-form score
In a first numerical experiment, we design a one-dimensional analytic toy model. Cubic spline data interpolation is used to build the surrogate $S(x)$ from the snapshots and the error bound is set to two times the absolute value of the true error $E(x)=2|S(x)-S^{\ast}(x)|$.
Using spline interpolation is not recommended in practice, as it easily creates wide oscillations that can artificially create spurious high score areas that mislead the sampling routines. It is nonetheless an interesting benchmark for difficult situations where the reduced modeling produces very sub-optimal approximations in some areas.
The score function, depicted in Figure 1, is defined for $x\in\mathcal{X}=\mathbb{R}_{+}$ by $S^{\ast}(x)=|\Psi^{\ast}(x)|$ with
$\Psi^{\ast}(x)=\left\{\begin{aligned} l_{\max}\quad&\textrm{if}\quad x\leq 1/l_{\max}\\ 1/x+f(x)\quad&\textrm{else}\end{aligned}\right.,$
where $f(x)=\alpha\left(\,{\rm l}{\rm l}_{x_{0}\leq x<x_{1}}\sin(x-x_{0})^{2}+\,{\rm l}{\rm l}_{x_{1}\leq x}(\sin(x_{1}-x_{0})^{2}-\beta(x-x_{1}))\right),$ given parameters $x_{1}>x_{0}>1/l_{\max}$ and $\alpha,\beta>0$. The reference distribution $\pi$ is a log-normal distribution such that $\ln(x)\sim\mathcal{N}(1.5,1.5)$. Note that here the rare event is $\{S^{*}\geq l_{\rm{max}}\}$. The parameter are set to $x_{0}=0.5$, $x_{1}=5$, $\alpha=15$, $\beta=0.1$ and $l_{\max}=90$, corresponding to the rare event probability (given in closed-form) and given in the table above.
@@ 4.1.2 Example #2: reduced basis approximation of a PDE score
In our second numerical experiment, the score function is defined as $S^{\ast}(x)=\|\Psi^{\ast}(x)\|_{p}$ with $p=1$ or $p=\infty$, $\Psi^{\ast}(x)$ being the high-fidelity numerical approximation of the solution of a bi-dimensional PDE parametrized by a $q$-dimensional vector $x=(x_{1}\cdots x_{q})^{\intercal}\in\mathcal{X}=\mathbb{R}_{+}^{q}$. The chosen PDE corresponds to the well-studied thermal block problem, which models heat diffusion in an heterogeneous media. The reference distribution for $x$ will be a log-normal distribution with independent components with $\ln(x_{i})\sim\mathcal{N}(1.5,1.5)$, for $i=1,\ldots,q$. The probability of interest will correspond to the occurrence of the temperature exceeding the given threshold $l_{\max}$ either in average (examples #1, #2a #2c) or locally (example #2b) over the spatial domain $\Omega=[0,1]^{2}$. PDE solutions will be approximated using the popular reduced basis (RB) method.
@@
where the function $\kappa(\cdot;x):\Omega\to\mathbb{R}_{+}$ represents a parametric diffusion coefficient function. This PDE models the stationary temperature field $\Psi$ of an heterogeneous medium which is uniformly heated in the domain and thermostatted at the boundary.
The high-fidelity approximation $\Psi^{\ast}(x)$ of the solution of this diffusion problem entails the solution of a variational problem. Precisely, let $\mathcal{V}$ denote the Hilbert space $\mathcal{H}^{1}_{0}(\Omega)$ and $\|\nu\|_{\mathcal{V}}=\|\nabla\nu\|_{L^{2}}$ the norm induced by the inner product defined over $\mathcal{V}$. For any $x\in\mathcal{X}$, we define the solution $\Psi^{\star}(x)\in\mathcal{V}$ of the variational problem $a(\Psi(x),\nu;x)=\int_{\Omega}\nu\mathrm{d}\Omega,\quad\forall\nu\in\mathcal{V},$ where $a:\mathcal{V}\times\mathcal{V}\to\mathbb{R}$ is the bilinear form $a(u,\nu;x)=\int_{\Omega}\kappa(\cdot;x)\nabla u\cdot\nabla\nu d\Omega$. We consider the piecewise constant diffusion coefficient $\kappa(\cdot;x)=\sum_{i=1}^{q}\,{\rm l}{\rm l}_{\Omega_{i}}(\cdot)x_{i}$ with the partition $\Omega=\cup_{i=1}^{q}\Omega_{i}$ and $\Omega_{i}$’s being non-overlapping squares. Setting the parameter space $\mathcal{X}=(\mathbb{R}^{*}_{+})^{q}$ implies that the bilinear form is strongly coercive, and the Lax-Milgram lemma ensures the variational problem has a unique solution.
@@
with the energy norm induced by the bilinear form $\|\nu\|=\sqrt{a(\nu,\nu;x)}$. The high-fidelity approximation $\Psi^{\ast}(x)$ is obtained in practice by solving a large system of $h$ linear equations. [22]. For the function $score$ defined as the $L^{p}$-norm, the rare event {$S^{\ast}(X)=\|\Psi^{\ast}(X)\|_{L^{p}}>l_{\max}\}$ corresponds to reaching a critical temperature $l_{\max}$, e.g. a melting temperature, on average in the domain ($p=1$) or in some point of the domain ($p=\infty$).
=$\mathcal{M}=\{\Psi^{\ast}(x)\in\mathcal{V}_{h}:x\in\mathcal{X}\},$
by evaluating only some of its elements. By selecting a well-chosen set of $K$ snapshots $\{\Psi^{\ast}(x_{i})\}_{i=1}^{K}$, the RB framework approximates $\mathcal{M}$ in a well-chosen linear subspace $\mathcal{V}_{K}\subset\mathcal{V}_{h}$ of much lower dimension $K\ll h$. Using this framework yields i) an approximation $\Psi$ of the high-fidelity solution $\Psi^{\ast}$ which is used to compute the approximation $S(x)=\|\Psi(x)\|_{L^{p}}$ of the score $S^{\ast}(x)$, ii) a so-called a posteriori error bound on $\|\Psi^{\ast}(x)-\Psi(x)\|_{\mathcal{V}_{h}}$ denoted by $\Delta_{\Psi}(x)$. The high-fidelity solution, the RB approximation and the a posteriori error bounds are computed via the pymor Python® toolbox [19] available at https://github.com/pymor/pymor.
To use the ARMS algorithm, we need to design an error bound $E(x)$ majoring the absolute error on the score $|S^{\ast}(x)-S(x)|$, relying on the a posteriori error bound on $\|\Psi^{\ast}(x)-\Psi(x)\|_{\mathcal{V}_{h}}$. In the case of $p=1$, the error bound satisfies
$|S^{\ast}(x)-S(x)|$ $=|\|\Psi^{\ast}(x)\|_{L^{1}}-\|\Psi(x)\|_{L^{1}}|$
$\leq\|\Psi^{\ast}(x)-\Psi(x)\|_{L^{1}}\leq\|\Psi^{\ast}(x)-\Psi(x)\|_{L^{2}}\leq\pi^{-1}\|\Psi^{\ast}(x)-\Psi(x)\|_{\mathcal{V}_{h}}$
$\leq\pi^{-1}\Delta_{\Psi}(x)=E(x),$
where we have used the triangular inequality, the Cauchy-Schwarz, the Poincaré inequality with the constante $\pi^{-1}$ on the unit square domain and the a posteriori error bound. As the finite element basis is an interpolating one, we will assume $h$ sufficiently large so that we can compute the $L^{1}$-norm as the average of the finite element coefficients. In the case of $p=\infty$, except in the one-dimensional case, one can not in general rigorously upper bound $\|\Psi^{\ast}(x)-\Psi(x)\|_{L^{\infty}}$ by $\|\Psi^{\ast}(x)-\Psi(x)\|_{\mathcal{V}_{h}}$. We will nevertheless use for this experiment the rough approximation $|S^{\ast}(x)-S(x)|\approx\|\Psi^{\ast}(x)-\Psi(x)\|_{\mathcal{V}_{h}}\leq\Delta_{\Psi}(x)=E(x)$, possibly violating assumption (1.2).
Let us comment on two important features of the RB approximation in regards to our rare event simulation framework. First, we emphasize that our algorithm enriches the RB approximation space using a generalization of the popular weak greedy algorithm. Indeed, the latter selects at each iteration the snapshot in the solution set $\mathcal{M}$ which possess the worst a posteriori error bound for the current RB approximation. This setup corresponding in our algorithm to the degenerated case where the learning parameter is set to $\tau=+\infty$. At the opposite, the setup where $\tau=0$ will correspond to sampling according to the proposal, independently of the approximation error. In the general setup where $0<\tau<\infty$, snapshot selection will balance these two cases, as the RB will be enriched with a sample drawn according to a weighted version of the proposal distribution, with weights proportional to the function $e^{\tau E^{(k)}}$.